Use the following probability distribution for this next question. Biologists researching a certain type of hawk, Accipiter spp., found the following probability values for x, the number of offspring.
x
P(x)
0
.05
1
.19
2
.32
3
.21
4
.12
5
.08
6
.03
Find the mean of the probability distribution for offspring of Accipiter spp. . Give your answer to one decimal, e.g., 1.2 .

Answer by Boreal(15235)

Mean=E(X)=sum of x*p(x)=0*0.05+1*0.19+2*0.32+3*0.21+4*0.12+5*0.08+6*0.03
=0+0.19+0.64+0.63+0.48+0.40+0.18
-2.52 or 2.5